Daily Reports

Because of COVID-19, only IRAM staff is allowed to observe on-site until new notice.

Back to the NIKA2 Pool page

Nov 15th, Monday

19:30-24:00, 10th IRAM 30m School - Observing slot of Working Group IV. Observers: Ioannis & Carsten. Aside from observing a nearby galaxy, we did several calib1 scans on Uranus and a Skydip. Weather conditions were good, stable (tau225~0.15).

Nov 16th, Tuesday

19:30-24:00, 10th IRAM 30m School - Observing slot of Working Groups II; Juan and Albrecht: observing OK with NIKA2 with tau225 < 0.2, stable; several maps on the Horsehead and W3OH; we got a calib_1scan on Uranus (s179) and a skydip (s181), project s02-21.

Nov 17th, Wednesday

00:35-03:00, Angel: Restarting acquisition, cryo run 68. Opacity below 0.2. Stable. Imbfits creation restarted twice: OK. Pointing/focus on W3OH. Three repetitions on PSZ2G141 (project 199-16).

A timelapse of Manuel Castillo Fraile from this morning:


03:00-09:00, Nicolas Ponthieu on shift. Starting with Angel and Sergio, telescope focused, pointing on Uranus, beam map (20211117s28), calib_1scan (s29), skydip (s30), then one rep of 199-16/PSZ2G141 to complete, then 192-16/cosmos until 9am with focus/pointing around 5h50 and 8h10. The night went smoothly.

09:00-15:00, Angel, Gianluca (invited for further observing training): Opacity around 0.2 about the first third of the shift. 192-16/cosmos continued for ~1h. Then, we got 1.5 reps. on AK4-224 cluster (084-21) with tau225 = 0.20 to 0.25. Some instabilities appeared after that, with opacity increasing from 0.3 to 0.4. Acquisition restarted at 48 Hz at 12h (UT; starting from scan 20211117s124) and 2 reps. were obtained on cluster A1835 of ARSENAL sample (086-21). A broadening of about 1-2 arcsec in the beam size with the best focus correction was observed at the last part of the shift. NIKA2 observations stopped after a calib_1scan on Mars (20211117s141; broadened beam) skydip (s143).

17:30-24:00, 10th IRAM 30m School - Working Group V (orange) observing on the Eyelash (project s05-21); got a calib_1scan on NGC7027 (20211117s232), CRL2688 (s233), MWC349 (s235) after pointing/focusing on the first source. Then, we did a beam map on 2251+158 (s244).

Nov 18th, Thursday

00:00-03:00, Angel: Starting the mini-shift with opacity about 0.2, stable. Two reps. done on pre-stellar core 1390 (007-21). Then, 2 reps. on NGC1333-C2 (010-21); <tau225> about 0.3 at the end.

03:00-09:00, Alexandre Beelen on shift, starting with Angel, telescope focused and pointed, one repetition on NGC1333-C2, pointing, 3 rep on NGC1333-C2, switched to 045-21, weather to variable tau not well fitted, pointing, bloc on GP212_3*, stop after 3 scans due to high wind, restarted, pointing 1 scan on GP212_3*, stop because of snow at 07h (LT).

09:00-15:00, Angel: Observations resumed at 12:40 (LT). Opacity about 0.25, but unstable sometimes according to the variations of pointing/focus baselines at 2 mm. More stable after 13:00. Two repetitions (= 4 scans) on NGC3938 (IMEGIN) before opacity increasing. Acquisition restarted at 48 Hz at 14:10 for changing to ARSENAL targets. Very unstable around 14:30, according to several pointing attempts. Stopped at 14:45, looking at the beginning of the first backup observing slot (with EMIR) on behalf of the 10th IRAM School.

15:00-00:00, 10th IRAM School - Observing slot of Working Group III (conducted by Mónica). Observing until 22:20. Telescope parked. Wind gusts (20-21 m/s) and light hail.

Nov 19th, Friday

Telescope stopped. Strong wind, hail.

Nov 20th, Saturday

Telescope stopped. Dense fog, snow.


Nov 21st, Sunday

00:00-09:00, telescope stopped. Foggy. Thick clouds.

09:00-15:00, Gianluca, Angel: the deicing system was switched off at ~08:30. Resuming telescope operations at 09:40. Crate 2 and 4 boards of Crate 2 were missing when the acquisition was started. Rebooting crates without effect. Fixed with a complete restart of the crates (power cycle). SN (standby) is aware. Acquisition restarted at 24 Hz, 0.3 < tau225 < 0.4 at the first attempt for pointing/focus on 1253-055. First focus corrections about +1 mm or a bit more, probably as an effect of temperature gradients in the quadrupod (spider). Focus corrections did move to values around zero in about 1 hour. In the meantime, instability and mean opacity values skyrocketed. Telescope stopped at 11:30. Vertex closed. Snowing again at ~12:30.

15:00-20:30, Suzanne and Lara on hold. Telescope stopped.

20:30-21:00, Angel: weather brings an opportunity, tau225 ~0.4, main dish and sub-reflector OK; starting acquisition at 45 Hz. Pointing/focus on 2037+511. Relatively stable. Started one repetition (=4 scans) on A2218 (086-21).

21:00-03:00, Antonio Ferragamo, Aurore Bacmann, and Aishwarya Paliwal: The opacity was extremely unstable and above tau225>0.4. Due to the bad tau225 conditions, we decided to observe the filler project 045-21 throughout the entire shift. As planned by Angel, at the beginning of the night we performed a skydip [20211121s67], then we made three calib_1scan on NGC7027 [20211121s80], CRL2688 [20211121s81], MWC34 [20211121s82].

Nov 22nd, Monday

03:00-04:10, Ankur: continuing observations of project 045-21, starting with the field GP233*. Opacity increased suddenly above 1.6. Dense fog. Cancelling observations. Telescope stopped at 04:10. It starts to snow at ~04:30.

04:30-03:00, Matteo, Gianluca, Stavroula, Lara, Marco, and Antonio on-hold. Telescope stopped due to snow and light hail. Taumeter readings apparently stuck between 12:16 and 19:30.

Nov 23rd, Tuesday

04:45, Angel: Opacity < 0.3 and wind gusts speed dropping to ~10 m/s. Checking the antenna. Sergio cleaned as much as possible the sub-reflector and quadrupod from icicles. Starting acquisition with a sampling frequency of 24 Hz. Pointing and focus on 0823+033, pretty stable (baseline variations +/- 1.5 Jy/beam (2 mm). Best focus is about +0.3 mm; a whole day of de-icing power cycles.

06:00-10:30, Ankur: After satisfactory pointing/focus, at 06:40 science observations started with project 085-21 and did 4 reps on source eFEDS3M. Then, did a couple of calib_1scan on Mars (s56, s64) and a skydip (s66). Tau ~ 0.12 all morning. Overall very stable conditions.

DailyReportsNika2Pool161121 (last edited 2021-11-23 10:07:00 by NikaBolometer)